# 1]FAIR at Meta 2]Univ. Grenoble Alpes, Inria, CNRS, Grenoble INP, LJK, France

### v1

@@ Abstract
[Website]johnlnguyen.com/oneflow
@@ 1 Introduction
Native Multimodal Models — models capable of handling both multimodal understanding and generation within a single backbone — have advanced considerably in visual understanding and generation. These models typically employ a unified transformer architecture with next-token prediction to handle both discrete and continuous generation (Team, 2024; Wu et al., 2025; Ma et al., 2025; Deng et al., 2025; Zhou et al., 2025). Recent work like Transfusion (Zhou et al., 2025) and Show-O (Xie et al., 2024) demonstrates that leveraging modality-specific training objectives within shared architectures can significantly improve performance, particularly on continuous modalities such as vision.
@@ 2.1 Discrete Text Generation via Edit Flows
2.1.2 Insertion prediction ($\lambda^{i}$).
@@
where $\pi\in(0,1)$ is the probability of observing zero, and $\lambda_{\text{nonzero}}>0$ is the rate parameter but restricted to only modeling the distribution of nonzero counts. We train $\pi$ by using a binary cross entropy (BCE) loss to detect if the missing count is zero, and we train $\lambda_{\text{nonzero}}$ using the original loss ((4)) on nonzero counts. For sampling (3), we can use the expectation $\lambda^{i}(X_{t})=(1-\pi^{i}(X_{t}))\lambda^{i}_{\text{nonzero}}(X_{t})$. However, we found that a consistently better sampling strategy is to first sample whether there are zero insertions using $\pi$, then simply use the rate $\lambda^{i}_{\text{nonzero}}(X_{t})$ if there are nonzero insertions. Pseudocode for sampling can be found in Algorithm 2.
2.1.3 Bag-of-tokens prediction ($Q^{i}$).
@@ 2.3 Concurrent Mixed-modal Generation
2.3.1 Independent mixed-modal generation.
@@
2.3.2 Interleaved mixed-modal generation.
@@ 3 Findings and Unlocking New Capabilities in Unified Multimodal
§3.43.5 How does OneFlow compare to other unified multimodal models?
@@
3.0.2 Datasets.
For multimodal pretraining, we use image-text pairs from a filtered version of the Conceptual Captions dataset (CC12M (Sharma et al., 2018)), the YFCC dataset (Thomee et al., 2016), and licensed data, for a total of 400M examples. During instruction finetuning, we use a filtered image portion of the PerceptionLM dataset (Cho et al., 2025), interleaving data from Chameleon (Team, 2024), and a filtered portion of the Cambrian-7M (Tong et al., 2024a) dataset.
3.0.3 Baselines.
To evaluate our model’s performance against existing methods, we compare against two baselines: (1) an autoregressive (AR) + Flow Matching (FM) multimodal model based on Transfusion (Zhou et al., 2025), where text tokens are generated autoregressively and image tokens via FM, and (2) a masked diffusion model based on LLaDA (Nie et al., 2025). For the masked diffusion baseline, we tested two sampling variants: low-confidence and random remasking, with random remasking performing better across all experiments. Unlike Transfusion, we follow Janus-Flow (Ma et al., 2025) and adopt a dual-encoder setup. For image encoders, we use a pretrained SigLIP2 ViT-SO400M-16@512 (Tschannen et al., 2025) for understanding and an SD3 VAE (Esser et al., 2024) for generation. Following Transfusion, we use U-Net adapters. Figure 19 illustrates how OneFlow consumes a mixed-modal sequence. To improve performance of the AR model, we follow Chameleon (Team, 2024) and sweep over several generation lengths for optimizing CIDEr score. For image generation, we use a first order Euler solver with entropy rectifying guidance (Ifriqi et al., 2025), we set the guidance scale to $5.0$ across $50$ sampling steps.
3.0.4 Model Configurations.
To investigate scaling trends, we train models at three different sizes – 1B, 3B, and 8B parameters, and token counts. For each benchmark, we plot all results over log-FLOPs curve and regress linear trendlines. Following Transfusion, we also estimate relative compute efficiency by measuring the parity FLOP ratio: the ratio between the number of FLOPs required by our AR and OneFlow to reach the same level of performance.
3.0.5 FLOPS Estimation.
To estimate the FLOPS during training, we track the number of tokens throughout training for all models. During multimodal pretraining, we do not delete images and OneFlow uses on average 50% fewer generated text tokens compared to AR. We use Flash Attention and follow their FLOP estimation (Dao et al., 2022).
3.0.6 Evaluation setup.
@@ 3.1 OneFlow Scales Better Than AR
In this experiment, we study the performance of OneFlow and AR in controlled settings at various model sizes and token counts. To ensure OneFlow has no advantage in data-constrained settings, we trained both models on 2B image-text pairs over 500k iterations using a batch size of 4096. Both models were initialized from Llama 3.2 1B (AI@Meta, 2024). For AR, the number of tokens predicted during training equals the sequence length, whereas for OneFlow, the number of predicted tokens corresponds to the number of deleted tokens, which on average is 50% of the data sequence. We measure compute efficiency using the parity FLOP ratio (Table 1), which quantifies the relative FLOPs required by OneFlow to match AR’s performance. Our results show that OneFlow achieves the largest compute savings on dense prompt alignment and captioning tasks, requiring less than half the FLOPs of AR to reach the same level.
We find that OneFlow scales better than AR on every benchmark. This scaling advantage is especially pronounced on DPG Bench, where OneFlow scales significantly better than AR. Conversely, for image captioning, OneFlow shows a notable performance gap relative to AR. Figure 5 visualizes the scaling trend, and the final metrics after training are shown in Table 2, along with a comparison against other state-of-the-art models.
@@
Parity FLOP Ratio 0.49 0.97 0.32 0.42
@@ 3.2 Mixed Modal Pre-training Enables Better Generation and Understanding
In this section, we study the impact of mixed modal pretraining. We investigate whether concurrent mixed modal pretraining and sequential pretraining affect downstream understanding tasks. We train two 1B variants: one using sequential pretraining (T2I or I2T), and another with the same data but where 20% of examples use concurrent generation.
Figure 6 indicates that when starting from a concurrent model, downstream image understanding improves by 5% relatively. These results demonstrate the impact of concurrent mixed modal pre-training dense prompt alignment. We also finetune the pretrained model from Section 3.1 on a mixture of interleaving data, filtered PLM data, and text data to evaluate downstream performance.
3.3 Hierarchical Generation Enables Reasoning
@@
In this section, we present the image generation results of our pretrained model compared to controlled baselines. We train both models on a mixture of image-text paired data. To ensure we do not give the OneFlow model any advantage in data-constrained settings, we train both models on 500M unique samples (roughly 256B tokens) and epoch over the data 4 times. We expect AR and OneFlow to behave similarly for image generation, since both use continuous flow matching (FM) for this task. Indeed, we observe similar FID scores across all three methods, and AR and OneFlow exhibit comparable scaling trends. The most substantial improvement comes from dense prompt alignment (DPG Bench), where OneFlow tends to capture subtle details more accurately than the AR model. We show qualitative examples in Figure 8.
@@
In this section, we present the results of the image understanding in the VQA and captioning tasks compared to the controlled baselines. We finetune both models on a large instruction tuning dataset totaling 40M samples. On average, OneFlow outperforms AR across all VQA tasks. The greatest improvement appears in vision-centric tasks such as RealWorldQA, where the bidirectional models (Mask and OneFlow) achieve 10% higher performance than AR.
General Knowledge OCR & Chart Vision Halluc. Model Params MMB VQAv2 GQA MME MMMU AI2D DocVQA ChartQA TextVQA RealWorld POPE Multimodal LM Show-O (Xie et al., 2024) 1.3B 61.0 1232.9 27.4 84.5 MetaMorph (Tong et al., 2024b) 7B 75.2 41.8 37.1 60.5 58.3 Janus-Flow (Ma et al., 2025) 1.5B 74.9 79.8 60.3 1333.1 29.3 64.6 55.5 88.0 Janus-Pro (Chen et al., 2025) 1.5B 73.4 67.9 59.3 1443.0 33.4 62.8 21.2 35.8 53.9 53.5 84.8 Janus-Pro (Chen et al., 2025) 7B 76.9 74.1 62.0 1531.0 38.2 68.1 24.3 57.2 56.4 85.2 Mask Diffusion Muddit (Shi et al., 2025) 1B 67.7 57.1 1104.6 D-DiT (Li et al., 2025) 2B 60.1 59.2 1124.7 84.0 MMADA (Yang et al., 2025) 8B 68.5 76.7 61.3 30.2 86.1 Discrete Flow FUDOKI (Wang et al., 2025) 1.5B 73.9 57.6 1485.4 34.3 86.1 Controlled Comparisons AR + FM Ablation 1B 64.7 66.0 55.0 1394.8 28.9 59.1 22.7 35.5 48.3 50.5 85.6 Mask + FM Ablation 1B 66.0 64.4 55.6 1462.2 28.4 55.1 18.7 34.7 44.6 50.6 84.6 OneFlow 1B 69.0 67.7 57.8 1497.1 29.8 58.5 23.8 35.0 50.4 50.6 85.7 OneFlow 8B 72.5 73.7 61.9 1542.5 33.1 63.4 37.1 42.1 58.6 54.0 86.3
@@ 3.5 Comparison with State-of-The-Art Unified Models
We compare OneFlow with other autoregressive and diffusion multimodal models and summarize the results in Table 3. OneFlow achieves competitive performance on understanding and generation benchmarks, matching state-of-the-art models despite differences in training procedures. For example, models such as MMaDA underwent extensive post-training, while OneFlow did not. Similarly, FUDOKI was initialized from a pretrained multimodal model, whereas ours was trained from scratch.
We evaluate FID using the open-source weights from these models on COCO-2014. We find that MMaDA and Bagel underwent extensive aesthetic finetuning that produces images stylistically distant from real images, resulting in higher FID scores. For a fair comparison, we evaluate using the open-source weights of MMaDA Base and Bagel without thinking mode for image generation.
@@ 3.6 OneFlow Unlocks New Capabilities
3.6.1 Classifier-free guidance improves text detailedness.
@@
The use of continuous-time Markov chains allows us to apply classifier-free guidance (CFG) to our model’s insertion rates. Specifically, given an unconditional prediction $\lambda(X_{t})Q(X_{t})$ and a conditional prediction $\lambda(X_{t}|c)Q(X_{t}|c)$, where $c$ is the prompt and $w$ is the guidance weight, the modified insertion rate is constructed as:
@@
3.6.2 Simultaneous generation of interleaved text and images.
@@ 4 Related Work
Figure 11: Illustration of the model input and targets during training for (top) text generation and (bottom) interleaved generation. To train autoregressive with diffusion denoising, the images are typically duplicated so that both the clean and the noisy images are in the sequence. On the other hand, OneFlow deletes tokens and images during training which reduces the sequence length.
Generative Model Framework Text AR AR Mask Diffusion Discrete FM Edit Flow Image AR Diffusion / FM Mask Diffusion Discrete FM FM Training Properties Attention Causal Block Causal Bidirectional Bidirectional Bidirectional Tokens / Iter Seq Len up to 2x Seq Len Seq Len Seq Len $\sim$50% Seq Len Generation Capabilities Understanding Single image gen Variable length $\times$ $\times$ Interleaved gen $\times$ $\times$ $\times$ Examples Chameleon (Team, 2024) JanusPro (Chen et al., 2025) Transfusion (Zhou et al., 2025) Bagel (Deng et al., 2025) MMaDA (Yang et al., 2025) FUDOKI (Wang et al., 2025) OneFlow
@@
Current approaches for unified multimodal models fall into three main paradigms: fully autoregressive (Team, 2024; Wang et al., 2024; Wu et al., 2025), hybrid (Zhou et al., 2025; Deng et al., 2025; Xie et al., 2024; Ma et al., 2025), and fully diffusion-based (Yang et al., 2025; Swerdlow et al., 2025; Li et al., 2025; Wang et al., 2025). While these models are limited by a fixed generation order or fixed-length output, our approach fundamentally differs by being able to simultaneously generate interleaved content and a variable number of images. We summarize the differences between each framework in Table 4.
@@
Iterative refinement models, including diffusion (Sohl-Dickstein et al., 2015; Ho et al., 2020) and flow models (Liu et al., 2022; Albergo et al., 2023; Lipman et al., 2024), have been adapted for discrete token spaces. Discrete diffusion models typically learn to reverse a corruption process (Austin et al., 2021; Lou et al., 2024), while discrete flow models transport between two distributions with an interpolating scheme (Campbell et al., 2024b; Gat et al., 2024). Although these frameworks offer a large design space (Shaul et al., 2024; Wang et al., 2025), recent works have predominantly focused on a simplified mask construction (Sahoo et al., 2024; Shi et al., 2024a; Ou et al., 2024; Zheng et al., 2024). This masking framework, however, cannot be easily applied to variable-length and especially simultaneous interleaved generation.
@@
Early non-autoregressive models for variable-length generation (Gu et al., 2019a, b; Stern et al., 2019; Reid et al., 2022) often relied on multiple models and evaluations to handle edit operations. While later work like Edit Flows (Havasi et al., 2025) improved on this by using a continuous-time framework and using only a single evaluation per step. Campbell et al. (2024a) also proposed modeling insertions with a diffusion model for denoising, but did not consider sequential data and requires a separate network due to using a single time during inference. In contrast, our approach considers mixed-modal sequence data, denoising images concurrently with per-image time values, allows for parallel token insertions, and uses a unified backbone architecture.

### v2

@@ Abstract
[Website]https://oneflow.framer.ai/
@@ 1 Introduction
Native Multimodal Models — models capable of handling both multimodal understanding and generation within a single backbone — have advanced considerably in visual understanding and generation. These models typically employ a unified transformer architecture with next-token prediction to handle both discrete and continuous generation (Team, 2024; Wu et al., 2025a; Ma et al., 2025a; Deng et al., 2025b; Zhou et al., 2025). Recent work like Transfusion (Zhou et al., 2025) and Bagel (Deng et al., 2025b) demonstrates that leveraging modality-specific training objectives within shared architectures can significantly improve performance, particularly on continuous modalities such as vision.
@@ 2.1.1 Parameterization
2.1.2 Insertion prediction $\lambda^{i}$
@@
where $\pi\in(0,1)$ is the probability of observing zero, and $\lambda_{\text{nonzero}}>0$ is the rate parameter but restricted to only modeling the distribution of nonzero counts. We train $\pi$ by using a binary cross entropy (BCE) loss to detect if the missing count is zero, and we train $\lambda_{\text{nonzero}}$ using the original loss ((4)) on nonzero counts. For sampling (3), we can use the expectation $\lambda^{i}(X_{t})=(1-\pi^{i}(X_{t}))\lambda^{i}_{\text{nonzero}}(X_{t})$. However, we found that a consistently better sampling strategy is to first sample whether there are zero insertions using $\pi$, then simply use the rate $\lambda^{i}_{\text{nonzero}}(X_{t})$ if there are nonzero insertions. We provide sampling pseudocode in Algorithm 2.
2.1.3 Bag-of-tokens prediction $Q^{i}$
@@ 2.3 Concurrent Mixed-modal Generation
2.3.1 Independent mixed-modal generation
@@
2.3.2 Interleaved mixed-modal generation
@@ 3 Findings and Unlocking New Capabilities in Unified Multimodal
§3.4 How does OneFlow compare to other unified multimodal models?
@@ 3.0.1 Training stages
3.0.2 Datasets
For multimodal pretraining, we use image-text pairs from a filtered version of the Conceptual Captions dataset (CC12M (Sharma et al., 2018)), the YFCC dataset (Thomee et al., 2016), and licensed data, for a total of 400M examples. During instruction finetuning, we use a filtered image portion of the PerceptionLM dataset (Cho et al., 2025), interleaving data from Chameleon (Team, 2024), and a filtered portion of the Cambrian-7M (Tong et al., 2024a) dataset, totaling around 40M examples.
3.0.3 Baselines
To evaluate our model’s performance against existing methods, we compare against two baselines: (1) an autoregressive (AR) + Flow Matching (FM) multimodal model based on Transfusion (Zhou et al., 2025), where text tokens are generated autoregressively and image tokens via FM, and (2) a masked diffusion model based on LLaDA (Nie et al., 2025). For the masked diffusion baseline, we tested two sampling variants: low-confidence and random remasking, with random remasking performing better across all experiments. Unlike Transfusion, we follow Janus-Flow (Ma et al., 2025a) and adopt a dual-encoder setup. For image encoders, we use a pretrained SigLIP2 ViT-SO400M-16@512 (Tschannen et al., 2025) for understanding and an SD3 VAE (Esser et al., 2024) for generation. Following Transfusion, we use U-Net adapters. Figure 20 illustrates how OneFlow consumes a mixed-modal sequence. To improve performance of the AR model, we follow Chameleon (Team, 2024) and sweep over several generation lengths for optimizing CIDEr score. For image generation, we use a first order Euler solver with entropy rectifying guidance (Ifriqi et al., 2025), we set the guidance scale to $5.0$ across $50$ sampling steps.
3.0.4 Model Configurations
To investigate scaling trends, we train models at three different sizes – 1B, 3B, and 8B parameters, and token counts. For each benchmark, we plot all results over log-FLOPS curve and regress linear trendlines. Following Transfusion, we also estimate relative compute efficiency by measuring the parity FLOPS ratio: the ratio between the number of FLOPS required by our AR and OneFlow to reach the same level of performance.
3.0.5 FLOPS Estimation
To estimate the FLOPS during training, we track the number of tokens throughout training for all models. During multimodal pretraining, we do not delete images and OneFlow uses on average 50% fewer generated text tokens compared to AR. We use Flash Attention and follow their FLOPS estimation (Dao et al., 2022).
3.0.6 Evaluation setup
@@ 3.1 OneFlow Scales Better than AR
We first discuss controlled experiments, where we study the performance across different frameworks when trained on the same data sets using the same backbone architecture. Specifically, we study the performance of OneFlow and AR in controlled settings at various model sizes and token counts. To ensure OneFlow has no advantage in data-constrained settings, we trained both models on 2B image-text pairs over 500k iterations using a batch size of 4096. Both models were initialized from Llama 3.2 1B (AI@Meta, 2024). For AR, the number of tokens predicted during training equals the sequence length, whereas for OneFlow, the number of predicted tokens corresponds to the number of deleted tokens, which on average is 50% of the data sequence. We measure compute efficiency using the parity FLOP ratio (Table 1), which quantifies the relative FLOPS required by OneFlow to match AR’s performance. Our results show that OneFlow achieves the largest compute savings on dense prompt alignment and captioning tasks, requiring less than half the FLOPS of AR to reach the same level.
We find that OneFlow scales better than AR on every benchmark. This scaling advantage is especially pronounced on DPG Bench, where OneFlow scales significantly better than AR. For image captioning as well, OneFlow shows a notable performance gap relative to AR. Figure 5 visualizes the scaling trend, and the final metrics are shown in Tables 2 and 3, along with a comparison against other state-of-the-art models. A detailed breakdown of the controlled comparisons among VQA task categories can be found in Table 17.
@@
Parity FLOPS Ratio 0.49 0.97 0.32 0.42
@@ 3.2 Mixed Modal Pre-training Enables Better Generation and Understanding
In this section, we study the impact of mixed-modal pretraining by comparing concurrent and sequential training approaches. Specifically, we train two 1B variants: one using sequential pretraining (either text-to-image or image-to-text), and another using the same data but with 20% of examples incorporating concurrent generation.
Our ablation experiment reveals that mixed-modal training significantly improves downstream visual question answering performance (Figure 6) and provides a minor boost to image generation. These results demonstrate the benefits of concurrent mixed-modal pretraining. Given these consistent improvements, we use the mixed-modal pretrained model for all subsequent finetuning experiments.
[…]
General Knowledge OCR & Chart Vision Halluc. Model Params MMB VQAv2 GQA MME MMMU AI2D DocVQA ChartQA TextVQA RealWorld POPE Multimodal LM Show-O (Xie et al., 2024) 1.3B 61.0 1232.9 27.4 84.5 MetaMorph (Tong et al., 2024b) 7B 75.2 41.8 37.1 60.5 58.3 Janus-Flow (Ma et al., 2025a) 1.5B 74.9 79.8 60.3 1333.1 29.3 64.6 55.5 88.0 Janus-Pro (Chen et al., 2025b) 1.5B 73.4 67.9 59.3 1443.0 33.4 62.8 21.2 35.8 53.9 53.5 84.8 Janus-Pro (Chen et al., 2025b) 7B 76.9 74.1 62.0 1531.0 38.2 68.1 24.3 57.2 56.4 85.2 Mask Diffusion Muddit (Shi et al., 2025) 1B 67.7 57.1 1104.6 D-DiT (Li et al., 2025) 2B 60.1 59.2 1124.7 84.0 MMaDA (Yang et al., 2025) 8B 68.5 76.7 61.3 30.2 86.1 Discrete Flow FUDOKI (Wang et al., 2025) 1.5B 73.9 57.6 1485.4 34.3 86.1 Controlled Comparisons AR + FM Ablation 1B 64.7 66.0 55.0 1394.8 28.9 59.1 22.7 35.5 48.3 50.5 85.6 Mask + FM Ablation 1B 66.0 64.4 55.6 1462.2 28.4 55.1 18.7 34.7 44.6 50.6 84.6 OneFlow 1B 69.0 67.7 57.8 1497.1 29.8 58.5 23.8 35.0 50.4 50.6 85.7 OneFlow 8B 72.5 73.7 61.9 1542.5 33.1 63.4 37.1 42.1 58.6 54.0 86.3
3.3 Hierarchical Generation Enables Emergent Reasoning
@@ 3.4 Comparison with State-of-The-Art Unified Models
In Tables 2 and 3, we benchmark OneFlow against other autoregressive and diffusion multimodal models. Our models deliver competitive performance across all tasks, rivaling state-of-the-art results despite significant differences in training procedures.
Understanding Performance.
On visual question answering, our 1B model nearly matches MMaDA 8B, while our 8B model surpasses MMaDA 8B on most benchmarks (Table 3). This is noteworthy because models like MMaDA received extensive post-training with chain-of-thought data and RL finetuning, whereas OneFlow did not. As discussed in Section 3.3, we demonstrate that reasoning naturally emerges during generation. Similarly, while FUDOKI leveraged a pretrained multimodal model, OneFlow delivers competitive performance without this advantage. We observe that initializing from a pretrained text-only autoregressive model consistently enhances both prompt alignment for image generation and VQA performance, yielding similar gains as the AR+FM model (Table 5).
Generation Performance.
For image generation, we measured FID using open-source weights on COCO-2014 (Lin et al., 2014) validation set. Since MMaDA and Bagel received extensive aesthetic finetuning that produces images stylistically distant from real images (yielding higher FID scores). For fair comparison, we benchmark against MMaDA Base and Bagel without thinking mode. We show qualitative comparisons in Figure 8.
@@ 3.5 OneFlow Unlocks New Capabilities
3.5.1 Classifier-free guidance improves text detailedness
@@
The use of continuous-time Markov chains allows us to apply classifier-free guidance (CFG) to our model’s insertion rates. Specifically, given an unconditional prediction $\lambda(X_{t})Q(X_{t})$ and a conditional prediction $\lambda(X_{t}|c)Q(X_{t}|c)$, where $c$ is the prompt and $w$ is the guidance weight, the modified insertion rate is:
@@
3.5.2 Simultaneous generation of interleaved text and images
@@ 4 Related Work
Figure 11: Illustration of the model input and targets during training for (top) text generation and (bottom) interleaved generation. On text generation, Edit Flow simply deletes tokens instead of replacing them with a special mask token, resulting in lower FLOPS but with the same information as the mask diffusion framework. On interleaved generation, to train autoregressive with diffusion denoising, the images are typically duplicated so that both the clean and the noisy images are in the sequence. On the other hand, OneFlow deletes tokens and images during training which reduces the sequence length.
Generative Model Framework Text AR AR Mask Diffusion Discrete FM Edit Flow Image AR Diffusion / FM Mask Diffusion Discrete FM FM Training Properties Attention Causal Block Causal Bidirectional Bidirectional Bidirectional Tokens / Iter Seq Len up to 2x Seq Len Seq Len Seq Len $\sim$50% Seq Len Generation Capabilities Understanding Single image generation Variable length $\times$ $\times$ Concurrent mixed-modal $\times$ $\times$ Interleaved generation $\times$ $\times$ $\times$ Examples Chameleon (Team, 2024) JanusPro (Chen et al., 2025b) Transfusion (Zhou et al., 2025) Bagel (Deng et al., 2025b) MMaDA (Yang et al., 2025) FUDOKI (Wang et al., 2025) OneFlow
@@
Early unified models like Flamingo (Alayrac et al., 2022) pioneered the integration of vision and language via cross-attention. Emu2 (Sun et al., 2023) advanced this paradigm by employing a unified autoregressive objective. Chameleon (Team, 2024) adopted an early-fusion, AR architecture that unifies images and text as discrete tokens. Subsequent work spans three architectural paradigms: diffusion (Yang et al., 2025; Li et al., 2025; Swerdlow et al., 2025; Shi et al., 2025; Wang et al., 2025), autoregressive (Alayrac et al., 2022; Team, 2024; Sun et al., 2024b; Chen et al., 2025b; Wu et al., 2025a; Yu et al., 2023), and hybrids (Zhao et al., 2024; Zhou et al., 2025; Xie et al., 2024; Tong et al., 2024b; Ma et al., 2025a; Shi et al., 2024b). Transfusion (Zhou et al., 2025) improves upon Chameleon by leveraging AR for text and diffusion for images. The Janus series collectively introduces decoupled visual pathways for understanding and generation (Wu et al., 2024; Ma et al., 2025b; Chen et al., 2025a). BAGEL scales multimodal pretraining using a Mixture‑of‑Transformers (MoT) (Liang et al., 2024; Deng et al., 2025a). In parallel, diffusion‑based models have emerged as promising alternatives: DDiT applies discrete diffusion to text and continuous diffusion to images (Li et al., 2025); UniDisc uses fully discrete diffusion with Llama tokenizer for text and MAGVIT‑v2 (Yu et al., 2024) for images (Swerdlow et al., 2025); FUDOKI uses discrete flow matching instead of masked diffusion, with a SigLIP (Zhai et al., 2023) encoder for understanding and a LlamaGen (Sun et al., 2024a) tokenizer for generation (Wang et al., 2025). MMaDA scales masked diffusion with mixed chain‑of‑thought fine‑tuning and UniGRPO (Yang et al., 2025). MuDDiT combines a pretrained T2I (Meissonic (Bai et al., 2024)) with a lightweight text decoder for MaskGIT‑style discrete diffusion (Chang et al., 2022; Shi et al., 2025). While these models are limited by a fixed generation order or fixed-length outputs, our approach fundamentally differs by simultaneously generating interleaved content and a variable number of images. We summarize the differences between each framework in Table 4.
@@
Iterative refinement models, including diffusion (Sohl-Dickstein et al., 2015; Ho et al., 2020) and flow models (Liu et al., 2022; Albergo et al., 2023; Lipman et al., 2024), have been adapted for discrete token spaces. Discrete diffusion models typically learn to reverse a corruption process (Austin et al., 2021; Lou et al., 2024), while discrete flow models transport between two distributions with an interpolating scheme (Campbell et al., 2024b; Gat et al., 2024). These frameworks offer a large design space for discrete generative modeling (Shaul et al., 2024; Wang et al., 2025), and can further be generalized to a large family of continuous-time Markov processes with a unified training recipe (Holderrieth et al., 2024). However, recent works on discrete diffusion have predominantly focused on a simplified mask construction (Sahoo et al., 2024; Shi et al., 2024a; Ou et al., 2024; Zheng et al., 2024). This masking framework, however, cannot be easily applied to variable-length and especially simultaneous interleaved generation.
@@
Early non-autoregressive models for variable-length generation (Gu et al., 2019a, b; Stern et al., 2019; Reid et al., 2022) often relied on multiple models and evaluations to handle edit operations, e.g. by inserting placeholder tokens first and then replacing them with the actual tokens. While later work like Edit Flows (Havasi et al., 2025) improved on this by using a continuous-time framework and removes the need for any placeholder tokens. In order to finetune from existing mask diffusion models, Wu et al. (2025b) and Kim et al. (2025) have also explored inserting only mask tokens, breaking down insertions back into two steps. In contrast, we simply made use of autoregressive models as initialization, modifying only model outputs for insertion predictions with no special tokens used in the model inputs, as this allows us to save on training FLOPS and inference costs. Related to combining discrete and continuous generative processes, Campbell et al. (2024a) also proposed modeling insertions with a diffusion model for denoising, but only considered singular insertions and required an additional generative model due to not using an interleaved time schedule. In contrast, our approach considers mixed-modal sequence data, denoising images concurrently with per-image time values, and allows for parallel token insertions.